OpenStudy (anonymous):

Use the quadratic formula to solve for x 8x^2 - 7x = -1 Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas.

OpenStudy (anonymous):

The answer I keep getting is \[x =\frac{ 7\pm \sqrt{17} }{ 16 }\]

OpenStudy (anonymous):

but my answer box will only let me type in individual answers...

OpenStudy (kidrah69):

ok we want all the numbers on one side so lets add one to both sides \[8x^2 - 7x +1=0\] now list the A B C values

OpenStudy (anonymous):

A = 8x^2 B = 7x C = 1

OpenStudy (kidrah69):

Good now plug them in \[x=\frac{ -b ± \sqrt(b^2 - 4ac) }{ 2a }\]

OpenStudy (kidrah69):

Wait

OpenStudy (kidrah69):

-7x=b

OpenStudy (kidrah69):

Plug it in a=8 b=-7 c=1

rvc (rvc):

or else find two numbers such that their \(\Large\ Product\) is 8 and their \(\Large\ Sum\) is -7

OpenStudy (kidrah69):

@rvc it says to use the quadratic formula

OpenStudy (kidrah69):

Not factor

rvc (rvc):

Ah okay :) but you can check this way too ;) plug in the values :)

OpenStudy (anonymous):

\[x = -7\pm \frac{ \sqrt{7^2-4*8*1} }{ 2*8 }\]

OpenStudy (anonymous):

ok sorry I had to retype out the equation like 3 times.. I kept messing up

OpenStudy (mathmate):

"If there is more than one solution, separate them with commas."

OpenStudy (kidrah69):

a neg times a neg is a pos :P \[\frac{-(-7) ± \sqrt((-7)^2 - 4(8)(1) }{ 2(8) }\]

OpenStudy (kidrah69):

You are mixing up your signs...

OpenStudy (mathmate):

Your initial answer is correct, which can be written as two answers: \(\frac{ 7+ \sqrt{17} }{ 16 }\),\(\frac{ 7 - \sqrt{17} }{ 16 }\)

OpenStudy (anonymous):

no I know I just can't seem to get to that point as far as I get is that quadratic equation but after that I'm not sure where to go to get the individual answers ( It won't let me do fractions or all of that though I literally have boxes to type in like real numbers but that's it)

OpenStudy (mathmate):

You may have to write it as (7+sqrt(17))/16,(7-sqrt(17))/16 but that depends on the application you're using. There would be help/instructions to guide you.

OpenStudy (mathmate):

@kidrah69 is right, "Round your answer to the nearest hundredth. " I missed that.

OpenStudy (mathmate):

So evaluate the numbers and type in two real numbers, separated by a comma. @Cassandra_Lea_96

OpenStudy (anonymous):

I don't know why that wasn't obvious to me... thanks aha

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

so 0.18 and 0.70 I changed the fractions to decimals then rounded to the nearest hundreth @mathmate

OpenStudy (mathmate):

The solutions are correct, and correctly rounded. Do not forget to separate the solutions by a comma as instructed. Instructions did not specify order, so you can answer in any order you want (i.e. 0.18,0.70 or 0.70,0.18 )

OpenStudy (anonymous):

ok =) thanks again!